# Stochastic Periodic Solutions for Newtonian Systems via Lyapunov Function

### v1

@@ 2 Preliminaries
$\lim_{R\to\infty}~{}\sup_{\|x\|>R,~{}t\in\mathbb{R}^{+}}\mathcal{L}\Psi(x,t)=-\infty,$ (2.4)
where the operator $\mathcal{L}$ is the generator of (2.1). Then (2.1) admits a $\mathcal{T}$-periodic Markov process.
It should be noted that there are several different ways of defining periodic solutions, but they are essentially the same. To the best of our knowledge, the earliest work on the periodic solution of stochastic processes can be traced back to Khasminskii [12, P. 64]. Ji et al. [10, 11] defined the stochastic periodic solution in terms of Fokker-Planck equations, while Feng et al. [5] defined it in terms of semigroups.
@@ 3 The Proof of Main Theorems
Here, $a>0$ is given in Hypothesis (H2) and related to convergent rate, and $D$ is a sufficiently large constant such that $\Psi\geq 1$ for any $(x,y,t)\in\mathbb{R}^{n}\times\mathbb{R}^{n}\times\mathbb{R}^{+}$. The first term is called to be the mixed energy. The selection of cross term between variables $x$ and $y$ is absolutely crucial when constructing Lyapunov functions for damped Newtonian systems. We choose the cross term to be $\langle y,\nabla F(x)-ax\rangle$ within the mixed energy, leading to significantly more concise computations and greatly advancing the applicable scope as described in Introduction.
@@
$\inf_{t\in\mathbb{R}^{+}}\Psi(x,y,t)\to+\infty,$
uniformly in time $t$ as $\|x\|+\|y\|$ tends to +$\infty$+ ∞. This implies function $\Psi$ satisfies condition (2.3) of Theorem 2.1.
The generator of system (1.6) is given by: for any $f\in C^{1,2,1}:\mathbb{R}^{n}\times\mathbb{R}^{n}\times\mathbb{R}^{+}\to\mathbb{R}$,
@@
$=$ $\partial_{t}f(x,y,t)+\frac{1}{2}\sum_{i,j=1}^{n}(\Sigma\Sigma^{\top})_{i,j}(x,y,t)\,\partial^{2}_{y_{i}\,y_{j}}f(x,y,t)$
@@
It remains to check condition (2.4) of Theorem 2.1. By computation, we have
@@
$\leq$ $\frac{1}{2}(c_{1}+c_{2}-1)\,(a\|y\|^{2}+b\|x\|^{2m})+\Big{(}M_{1}+\frac{1}{2}M_{2}+M+M^{*}\Big{)},$ (3.7)
for some positive constant $M^{*}$, where the first inequality (3.6) has used Hypotheses (H5), (H3) and $\|E(x,y,t)\|\leq e$, while the second inequality (3.7) has used Hypothesis (H4). This demonstrates (2.4) from the fact that $10<c_{1}+c_{2}<$ 1 and then the conclusion follows immediately from Theorem 2.1. $\Box$
Proof of Theorem 1.3.  The proof of this theorem is mainly obtained by applying Theorem 1.2, so we only need to verify Hypotheses (H1)-(H5) one by one. The proof is divided into several steps. It is easy to see that Hypothesis (H1) is satisfied by our assumption here.
@@
$\Big{[}V(x)+aF(x)-\frac{a^{2}}{2}\|x\|^{2}\Big{]}\geq\Big{(}\lambda+\lambda a-\frac{a^{2}}{2}\Big{)}\,\|x\|^{2},$
@@
[…]
@@
Thus far, we have verified Hypotheses (H1)-(H4). It is straightforward to verify that Hypothesis (H5) holds for $m=2(p+q-1)$.
@@
This implies the function $\Psi$ satisfies condition (2.3) of Theorem 2.1. In addition, using Hypotheses (A2)-(A4), we have
@@
which implies the function $\Psi$ satisfies condition (2.4). The conclusion follows immediately from Theorem 2.1. $\Box$
@@ 4 Plasma Physics Models with Bounded Potential Force
We claim that if $\|\Sigma\|_{2}^{2}\leq c_{3}(\|x\|^{2}+\|y\|^{2})+C$ with $c_{3}<16$, then it admits a $2\pi$-periodic solution in distribution.
@@
$-\langle\nabla_{x}V(x,t),\nabla F(x)-8x\rangle=-\frac{4\|x\|^{2}\,(2\|x\|^{2}-3)}{2+\sin t+\|x\|^{2}}~{}\leq-8\|x\|^{2}+36.$ (4.1)
It is easy to see that Hypotheses (H1)-(H5) hold with $b=8,\,m=1,$ and $M=36$. $\Box$
@@
We claim that if $\|\Sigma\|_{2}^{2}\leq c_{4}(\|x\|^{3}+\|y\|^{2})+C$ with $c_{4}<4$, then it admits a $2\pi$-periodic solution in distribution.
@@
$-\langle\nabla_{x}V(x,t),\nabla F(x)-2x\rangle=-\frac{4\|x\|^{4}}{\sqrt{2+\sin t+\|x\|^{2}}}\leq-2\|x\|^{3}+2.$ (4.2)
Hence, Hypotheses (H1)-(H5) hold with $b=2,\,m=3/2,$ and $M=2$. Then we conclude from Theorem 1.2 that there exists a $2\pi$-periodic solution in distribution. $\Box$
@@
[…]
We claim that if $\|\Sigma\|_{2}^{2}\leq c_{5}(\|x\|^{4}+\|y\|^{2})+C$ with $c_{5}<2$, then it admits a $2\pi$-periodic solution in distribution.
@@
Obviously, Hypotheses (H1)-(H5) hold with $b=1,\,m=2,$ and $M=3$. Then it yields from Theorem 1.2 that there exists a $2\pi$-periodic solution in distribution. $\Box$
@@
[…]
@@ 5 Conclusions and Further Perspectives
This paper proves the existence of a distributed periodic solution for stochastic time-periodic Newton systems under the assumptions that both the friction and potential functions tend to positive infinity at infinity, and the friction-potential gradient inner product grows at least like an even-power polynomial. This result is robust regarding bounded perturbations of the Hessian-friction matrix, at the cost of a slight increase in the growth rate of the potential function at infinity. Consequently, our method can be applied to systems like the one in Eq. (1.6), where $V(x)=\|x\|^{6}$ and $D^{2}F(x)$ is replaced by $D^{2}\left(\|x\|^{4}\pm\|x\|^{2}\right)+B(x,y,t)$ with $B$ being bounded. To the best of our knowledge, none of the existing results can be applied to this system. Our results not only largely confirm the stochastic Levinson conjecture proposed in [9] and resolve the open problem and the interesting problem posed in [9, P. 342], but also possess advantages 1-6 enumerated in the Introduction, which are absent in existing results.

### v2

@@ 2 Preliminaries
$\lim_{\|z\|\to\infty}~\mathcal{L}\Psi(z,t)=-\infty,\ {\rm uniformly\ in}\ t\in\mathbb{R},$ (2.7)
then (2.2) admits a periodic probability solution.
Note that (2.6) and (2.7) obviously imply (2.4) and (2.5), respectively. Under the assumption of locally Lipschitz continuity of coefficients, Corollary 2.4 is [12, Theorem 3.8].
From now on we assume that the drift $b:\mathbb{R}^{n}\times\mathbb{R}\to\mathbb{R}^{n}$ and noise intensity $\sigma:\mathbb{R}^{n}\times\mathbb{R}\to\mathbb{R}^{n\times n}$ are continuous on $\mathbb{R}^{n}\times\mathbb{R}$, locally Lipschitz continuous in their first variable $z$ and $\mathcal{T}$-periodic in their second variable $t$. Under this assumption, the transition probabilities associated to solutions of (2.1) can be defined. Let $Z^{s,z}(t)$ be the solution of (2.1) and ${\mathbb{P}}(s,z,t,A):=\mathbb{P}\big{(}Z^{s,z}(t)\in A\big{)}$ its transition probability function. Then ${\mathbb{P}}(s,z,t,A)$ is $\mathcal{T}$-periodic in the sense that ${\mathbb{P}}(s+\mathcal{T},z,t+\mathcal{T},\cdot)\equiv{\mathbb{P}}(s,z,t,\cdot)$ for all $z\in\mathbb{R}^{n}$ and $s\leq t$, see [12, Theorem 3.4]. We call a Markov process $Z(t)$ to be $\mathcal{T}$-periodic if its law ${\mathbb{P}}_{0}(t,A):={\mathbb{P}}\{Z(t)\in A\}$ satisfies the equation
${\mathbb{P}}_{0}(s,A)=\int_{\mathbb{R}^{n}}{\mathbb{P}}_{0}(s,dz)\,{\mathbb{P}}(s,z,s+\mathcal{T},A)\equiv{\mathbb{P}}_{0}(s+\mathcal{T},A),\quad\forall s\in\mathbb{R},\ A\in\mathcal{B}(\mathbb{R}^{n}),$ (2.8)
see [12, P. 64,  (3.17)]. (2.8) is equivalent to that the laws of $Z(s)$ and $Z(s+\mathcal{T})$ are identical for any $s\in\mathbb{R}$. According to [8, 9], such a $\mathcal{T}$-periodic solution $Z(t)$ is called to be periodic solution in distribution.
Ji et al [10, p.3-4] commented that the distribution of solutions to (2.1) is governed by (2.2) at least when the coefficients are locally Lipschitz continuous. Therefore, the distribution of periodic solutions to (2.1) correspond to periodic probability solutions to (2.2). We will give a proof for such a statement and further discuss the relations between the concepts of periodic probability solutions and periodic solutions in distribution.
Suppose that $(Z(t))_{t\in\mathbb{R}}$ is a solution of (2.1) whose law or distribution is $\mathcal{T}$-periodic, i.e., let $\mu_{t}=\mathbb{P}\circ(Z(t))^{-1}$, then $\mu_{t}=\mu_{t+\mathcal{T}},\ \forall t\in\mathbb{R}$. We claim that (2.3) holds.
For a given $\phi\in C^{2,1}_{0}(\mathbb{R}^{n}\times\mathbb{R})$, we can find a positive integer $k$ such that ${\rm supp}(\phi)\subset\overline{B_{k}(O)}\times[-k\mathcal{T},k\mathcal{T}]$. An application to $\phi$ of Itô’s formula yields
$\begin{split}\mathbb{E}_{\mu_{s}}\phi(Z(t),t)-\int_{\mathbb{R}^{n}}\phi(\cdot,s){\rm d}\mu_{s}&=\int_{s}^{t}\mathbb{E}_{\mu_{s}}\mathcal{L}\phi(Z(r),r){\rm d}r\\ &=\int_{s}^{t}\int_{\mathbb{R}^{n}}\mathcal{L}\phi(z,r){\rm d}\mu_{r}{\rm d}r,\quad s\leq t.\end{split}$
Let $s<-k\mathcal{T}$ and $t>k\mathcal{T}$. It follows from ${\rm supp}(\phi)\subset\overline{B_{k}(O)}\times[-k\mathcal{T},k\mathcal{T}]$ and the above equality that
$0=\int_{s}^{t}\int_{\mathbb{R}^{n}}\mathcal{L}\phi(z,r){\rm d}\mu_{r}{\rm d}r=\int_{\mathbb{R}}\int_{\mathbb{R}^{n}}\mathcal{L}\phi(z,r){\rm d}\mu_{r}{\rm d}r.$
This proves the claim, i.e., $(\mu_{t})_{t\in\mathbb{R}}$ is a periodic probability solution of (2.2) according to [10, 11]. This indicates that their periodic probability solution theory generalizes the classical periodic solutions of SDEs and accommodates coefficients with merely Sobolev regularity.
Conversely, assume that $(\mu_{t})_{t\in\mathbb{R}}$ is a periodic probability solution of (2.2) according to [10], let $Z^{s}(t),\ s\leq t$ be a solution of (2.1) with initial distribution $\mu_{s}$ and $\tilde{\mu}^{s}_{t}$ the distribution of $Z^{s}(t)$. In particular, $\tilde{\mu}^{s}_{s}=\mu_{s}$. Then we have
$\tilde{\mu}_{s+\mathcal{T}}(A)=\int_{\mathbb{R}^{n}}\mu_{s}(dz)\,{\mathbb{P}}(s,z,s+\mathcal{T},A),\quad\forall\ A\in\mathcal{B}(\mathbb{R}^{n}).$ (2.9)
An application of Itô’s formula yields
$\mathbb{E}_{\mu_{s}}\phi(Z^{s}(t),t)-\int_{\mathbb{R}^{n}}\phi(\cdot,s){\rm d}\mu_{s}=\int_{s}^{t}\mathbb{E}_{\mu_{s}}\mathcal{L}\phi(Z^{s}(r),r){\rm d}r=\int_{s}^{t}\int_{\mathbb{R}^{n}}\mathcal{L}\phi{\rm d}\tilde{\mu}^{s}_{r}{\rm d}r,\quad\forall\ \phi\in C^{2,1}(\mathbb{R}^{d}\times\mathbb{R}).$
In other words,
$\int_{\mathbb{R}^{n}}\phi(\cdot,t){\rm d}\tilde{\mu}^{s}_{t}=\int_{\mathbb{R}^{n}}\phi(\cdot,s){\rm d}\tilde{\mu}^{s}_{s}+\int_{s}^{t}\int_{\mathbb{R}^{n}}\mathcal{L}\phi{\rm d}\tilde{\mu}^{s}_{r}{\rm d}r,\quad\forall\ s\leq t,\ \phi\in C^{2,1}(\mathbb{R}^{d}\times\mathbb{R}).$ (2.10)
That is, $\tilde{\mu}^{s}_{t}$ satisfies (2.2) with the initial distribution $\mu_{s}$ in the sense of distribution. Note that [11, Lemma 2.1 (1)] ensures that $\{\tilde{\mu}^{s}_{t}\}_{t\geq s}$ is also a global probability solution of (2.2) with the initial distribution $\mu_{s}$. Suppose that the Cauchy problem associated with (2.2) admits a unique global probability solution. Then $\mu_{t}=\tilde{\mu}^{s}_{t}=\mathbb{P}[Z^{s}(t)\in\bullet]$ for all $t\geq s$. This shows that
$\mu_{s+\mathcal{T}}(A)=\int_{\mathbb{R}^{n}}\mu_{s}(dz)\,{\mathbb{P}}(s,z,s+\mathcal{T},A)=\mu_{s}(A),\quad\forall\ s\in\mathbb{R},\ A\in\mathcal{B}(\mathbb{R}^{n}).$
This illustrates the following: under the assumptions that the coefficients are locally Lipschitz continuous and that the probabilistic solution for Cauchy problem of system (2.2) possesses global uniqueness, the concept of a periodic probability solution in [10, 11] coincides with the concept of a periodic solution in distribution in [8, 9]. Under the non-degenerate assumption, [11, Corollary A] guarantees that such a uniqueness of global probability solution. However, stochastic Newtionian systems (1.10) are always degenerate. Fortunately, in many cases, it can be verified that (2.10) satisfies the Hörmander condition so that the Fokker-Planck operator becomes hypoelliptic. Given the similarities between hypoelliptic and elliptic operators, it seems promising to prove the uniqueness of global probability solutions, under additional smooth assumption of coefficients and Lyapunov dissipative condition. We leave it for future study.
Based on the above discussion, we adopt the more broadly applicable periodic probability solution as the definition for distributed periodic solution throught this paper.
The following example is taken from [9, subsection 3.1]. They made many numerical experiments about it. Their results shows the solutions starting from (0, 0) do not admit periodicity pathwisely. However, the sample paths do admit periodic phenomenon significantly in distribution. These phenomena confirm them that it is natural and essential to study stochastic dynamical problems in the sense of distribution. In the following, we explicitly solve its density of the $2\pi$-periodic probability solution or $2\pi$-periodic solution in distribution $(\mu_{t})_{t\in\mathbb{R}}$ for (2.11), which helps readers to understand these concepts.
Example 1
Consider the scalar periodic stochastic Duffing equation:
$\ddot{x}+\dot{x}+2x=\dot{B}_{t}+\cos t.$
Equivalently, we have
$\begin{cases}dx_{t}=y_{t}\,dt,\\ dy_{t}=(-2x_{t}-y_{t}+\cos t)\,dt+dB_{t}.\end{cases}$ (2.11)
Then define $Z:=(x,y)^{\top}$, it can be rewritten as
$dZ_{t}=(AZ_{t}+b(t))\,dt+\Sigma\,dB_{t},$
where
$A=\left(\begin{matrix}0&1\\ -2&-1\end{matrix}\right),\quad b(t)=\left(\begin{matrix}0\\ \cos t\end{matrix}\right),\quad\Sigma=\left(\begin{matrix}0\\ 1\end{matrix}\right)\quad~\hbox{and}\quad~\Sigma\Sigma^{\top}=\left(\begin{matrix}0&0\\ 0&1\end{matrix}\right).$
Its Fokker-Planck equation is
$\frac{\partial p(x,y,t)}{\partial t}=\frac{1}{2}\frac{\partial^{2}p(x,y,t)}{\partial{y}^{2}}-y\frac{\partial p(x,y,t)}{\partial x}+(2x+y-\cos t)\frac{\partial p(x,y,t)}{\partial y}+p(x,y,t).$ (2.12)
Explicitly, we can obtain
$p(x,y,t)=\frac{\sqrt{2}}{\pi}\,\exp{\left(-2\left(x-\frac{\cos t+\sin t}{2}\right)^{2}-\left(y-\frac{\cos t-\sin t}{2}\right)^{2}\right)},$
which is the density function of $2\pi$-periodic probability solution $(\mu_{t})_{t\in\mathbb{R}}$ for system (2.11). $\Box$
@@ 3 The Proof of Main Theorems
Here, $a>0$ is given in Hypothesis (H2) and related to convergent rate, and $D$ is a sufficiently large constant such that $\Psi\geq 1$ for any $(x,y,t)\in\mathbb{R}^{n}\times\mathbb{R}^{n}\times\mathbb{R}$. The first term is called to be the mixed energy. The selection of cross term between variables $x$ and $y$ is absolutely crucial when constructing Lyapunov functions for damped Newtonian systems. We choose the cross term to be $\langle y,\nabla F(x)-ax\rangle$ within the mixed energy, leading to significantly more concise computations and greatly advancing the applicable scope as described in Introduction.
@@
$\lim_{\|x\|+\|y\|\rightarrow+\infty}\Psi(x,y,t)=+\infty$
uniformly in time $t\in\mathbb{R}$. This implies function $\Psi$ satisfies condition (2.6) of Corollary 2.4.
The generator of system (1.6) is given by: for any $f\in C^{1,2,1}:\mathbb{R}^{n}\times\mathbb{R}^{n}\times\mathbb{R}\to\mathbb{R}$,
@@
$=$ $\partial_{t}f(x,y,t)+\frac{1}{2}\sum_{i,j=1}^{n}(\Sigma\Sigma^{\top})^{i,j}(x,y,t)\,\partial^{2}_{y_{i}\,y_{j}}f(x,y,t)$
@@
It remains to check condition (2.7) of Corollary 2.4. By computation, we have
@@
$\leq$ $\frac{1}{2}(c_{1}+c_{2}-1)\,(a\|y\|^{2}+b\|x\|^{2m})+\Big{(}M_{1}+\frac{1}{2}M_{2}+M+M^{*}\Big{)}$ (3.7)
for some positive constant $M^{*}$, where the first inequality (3.6) has used Hypotheses (H5), (H3) and $\|E(x,y,t)\|\leq e$, while the second inequality (3.7) has used Hypothesis (H4). This demonstrates (2.7) from the fact that $0<c_{1}+c_{2}<1$ and then the conclusion follows immediately from Corollary 2.4. $\Box$
Proof of Theorem 1.3.  The proof of this theorem is mainly obtained by applying Theorem 1.2, so we only need to verify Hypotheses (H1)-(H5) one by one. The proof is divided into several steps. It is easy to see that Hypothesis (H1) is satisfied by our assumptions here.
@@
$\Big{[}V(x)+aF(x)-\frac{a^{2}}{2}\|x\|^{2}\Big{]}\geq\Big{(}\lambda+\lambda a-\frac{a^{2}}{2}\Big{)}\,\|x\|^{2}$
@@
$=$ $2p\,\|x\|^{2(p-1)}\langle x,\nabla Q_{2q}(x)\rangle+\hbox{ lower-order\ terms}$
$=$ $4pq\,\|x\|^{2(p-1)}Q_{2q}(x)+\hbox{ lower-order\ terms}$
@@
Thus far, we have verified Hypotheses (H1)-(H4). It is straightforward to verify that Hypothesis (H5) holds for m=2(p+q1)m=2(p+q-1).
@@
This implies the function $\Psi$ satisfies condition (2.6) of Corollary 2.4. In addition, using Hypotheses (A2)-(A4), we have
@@
which implies the function $\Psi$ satisfies condition (2.7). The conclusion follows immediately from Corollary 2.4. $\Box$
@@ 4 Plasma Physics Models with Bounded Potential Force
We claim that if $\|\Sigma\|_{2}^{2}\leq c_{3}(\|x\|^{2}+\|y\|^{2})+C$ with $c_{3}<16$, then it admits a $2\pi$-periodic probability solution.
@@
$\langle\nabla_{x}V(x,t),\nabla F(x)-8x\rangle=-\frac{4\|x\|^{2}\,(2\|x\|^{2}-3)}{2+\sin t+\|x\|^{2}}~\leq-8\|x\|^{2}+36.$ (4.1)
It is easy to see that Hypotheses (H1)-(H5) hold with b=8,m=1,b=8,\,m=1, and $M=36$. $\Box$
@@
We claim that if $\|\Sigma\|_{2}^{2}\leq c_{4}(\|x\|^{3}+\|y\|^{2})+C$ with $c_{4}<4$, then it admits a $2\pi$-periodic probability solution.
@@
$\langle\nabla_{x}V(x,t),\nabla F(x)-2x\rangle=-\frac{4\|x\|^{4}}{\sqrt{2+\sin t+\|x\|^{2}}}\leq-2\|x\|^{3}+2.$ (4.2)
Hence, Hypotheses (H1)-(H5) hold with b=2,m=3/2,b=2,\,m=3/2, and $M=2$. Then we conclude from Theorem 1.2 that there exists a $2\pi$-periodic probability solution. $\Box$
@@
$V(x,t)=(2+\sin t)(1-\exp{(-\|x\|^{2})}),\qquad F(x)=\frac{1}{4}\int_{0}^{\|x\|^{2}}e^{s}\,(s+1)\,ds.$
We claim that if $\|\Sigma\|_{2}^{2}\leq c_{5}(\|x\|^{4}+\|y\|^{2})+C$ with $c_{5}<2$, then it admits a $2\pi$-periodic probability solution.
@@
Obviously, Hypotheses (H1)-(H5) hold with b=1,m=2,b=1,\,m=2, and $M=3$. Then it yields from Theorem 1.2 that there exists a $2\pi$-periodic probability solution. $\Box$
@@
In Examples 2-4, the same conclusions hold by Theorem 1.2 if the friction functions are replaced by $F(x)=\|x\|^{4}-|x\|^{2}$ and $F(x)=\frac{1}{4}\int_{0}^{\|x\|^{2}}e^{s}\,(s-1)\,ds$, respectively. Meanwhile, these correspond to stochastic nonlinear relaxation oscillation systems.
@@ 5 Conclusions and Further Perspectives
This paper proves the existence of a distributed periodic solution for stochastic time-periodic Newton systems under the assumptions that both the friction and potential functions tend to positive infinity at infinity, and the friction-potential gradient inner product grows at least like an even-power polynomial. This result is robust regarding bounded perturbations of the Hessian-friction matrix, at the cost of a slight increase in the growth rate of the potential function at infinity. Consequently, our method can be applied to systems like the one in Eq. (1.6), where $V(x)=\|x\|^{6}$ and $D^{2}F(x)$ is replaced by $D^{2}\left(\|x\|^{4}\pm\|x\|^{2}\right)+B(x,y,t)$ with $B$ being bounded. To the best of our knowledge, none of the existing results can be applied to this system. Our results not only largely confirm the stochastic Levinson conjecture proposed in [9] under broad conditions and resolve the open problem and the interesting problem posed in [9, P. 342], but also possess advantages 1-6 enumerated in the Introduction, which are absent in existing results.